OpenStudy (anonymous):

Please help for Medal!! What is the slope of the line that passes through the points (3, –1) and (–2, –5)?

OpenStudy (anonymous):

Do u hav choices?

OpenStudy (anonymous):

Assuming the points are (0,2) and (5,0) and that the "-" is used as a separator rather than as a minus sign. If it is a minus sign, why -0? Slope = difference in y coordinates/difference in x coordinates = (2 - 0)/(0 - 5) = 2/(-5) = -0.4

OpenStudy (anonymous):

A. -4/5 B.4/5 C.5/4 D.-5/4